OpenStudy (anonymous):

The height of a triangle is five more than three times the length of its base, . Which expression represents the area of the triangle?

OpenStudy (anonymous):

please help been stuck on this problem for almost an hour

OpenStudy (anonymous):

Since the area of a triangle is equal to 1/2base*height we can substitute the values into this equation. Let's assign x to the base. Therefore: Height = 5 + 3x so... Area = 1/2 *x(5+3x) and using the distributive property of multiplication... Area = 1/2 * (5x + 3x^2)